# FPBoost: Fully Parametric Gradient Boosting for Survival Analysis

### v1

@@ 3 Fully Parametric Survival Boosting
where $\alpha\geq 0$ is a weighting parameter and $\gamma\in[0,1]$ controls the ratio between the $L1$ and $L2$ penalties.
The training is performed using a standard gradient boosting algorithm. Specifically, for each parameter $\theta_{j}\in\Theta$, an empty list of trees is initialized. Then, at each iteration $m$, each of these lists is populated by a new tree $\tau_{j}^{(m)}(\mathbf{x})$ fitted on the negative gradient of the loss function, called pseudo-residual. Parameter estimation occurs by summing the estimations of all trees belonging to each list, weighted by a learning rate $\lambda>0$ [7]. The pseudocode of FPBoost is provided in Algorithm 1.
While the idea of leveraging a mixture of parametric functions is not new in the literature [17], nor is ensemble learning [13, 22], FPBoost proposes several innovations. First, the weighted sum is applied directly to the hazard function, unlike previous works that applied it to the survival or cumulative hazard function. Given the prominence and popularity of models based on hazard assumptions, such as proportionality or accelerated failure time, we argue that directly learning parameters in hazard-space can be beneficial for effective training. Second, FPBoost directly maximizes the survival likelihood, without relying on simplified custom losses such as the partial likelihood of proportional models or discrete losses of neural-network-based models. This is possible due to the assumption that the hazard is a composition of differentiable parametric hazards. Lastly, FPBoost leverages decision tree ensembles for parameter estimation, which have been historically extremely effective in dealing with tabular data, even against neural networks and deep learning [9].
To the best of our knowledge, FPBoost is the first model to combine tree-based learning via gradient boosting, composition of fully parametric hazards, and direct optimization of the survival likelihood into an empirically competitive survival analysis architecture.
@@
0:  Training data $\{(\mathbf{x}_{i},\delta_{i},t_{i})\}_{i=1}^{N}$, number of heads $J$, distribution types for each head, number of iterations $M$, learning rate $\lambda$, regularization parameters $\alpha$ and $\gamma$.
1:  Initialize $3J$ parameter models $F^{(0)}_{j}(\mathbf{x})$ with random values.
@@
3:     Compute $\Theta$ from the parameter models such that its $j$-th entry is $\theta_{j}=F^{(m)}_{j}(\mathbf{x})$.
4:     Apply a ReLU activation to $\Theta$.
5:     $L_{\text{lik}}=-\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{i}\log(h(t_{i}|\Theta))-H(t_{i}|\Theta)$
6:     $L_{\text{reg}}=\alpha\left(\gamma\left\|\Theta\right\|_{1}+(1-\gamma)\left\|\Theta\right\|_{2}^{2}\right)$
7:     $L=L_{\text{lik}}+L_{\text{reg}}$
8:     for each $\theta_{j}\in\Theta$ do
9:        Compute the pseudo-residual $r_{j}=-\partial L/\partial\theta_{j}|_{\Theta}$.
10:        Fit a regression tree $\tau_{j}^{(m)}(\mathbf{x})$ to $r_{j}$.
11:        Update the parameter model with the regression tree as $F^{(m+1)}_{j}(\mathbf{x})=F^{(m)}_{j}(\mathbf{x})+\lambda\tau_{j}^{(m)}(\mathbf{x})$.
12:     end for
13:  end for
14:  return  Trained FPBoost parameter models $F^{(M)}_{j}(\mathbf{x})$.
@@ 4.1 Datasets
To ensure fair evaluation and consistency with similar studies, we selected datasets from well-known benchmarks in survival analysis, covering different conditions like breast cancer, lung cancer, AIDS, and cardiovascular diseases. Table 2 collects the summary statistics of these datasets.
The AIDS [23] dataset originates from a trial comparing three-drug and two-drug regimens in HIV-infected patients. The primary event of interest was the time to an AIDS-defining event or death. The high censoring percentage resulted from the trial being terminated early after reaching a predefined level of statistical significance.
The Breast Cancer [24] dataset is derived from a study aimed at validating a 76-gene prognostic signature for predicting distant metastases breast cancer patients. The study includes gene expression profiling of frozen samples from 198 patients.
The German Breast Cancer Study Group (GBSG2) [25] dataset targets breast cancer recurrence post-treatment, evaluating hormone therapy’s impact on recurrence. Collected from a randomized study in Germany, it includes covariates like age, menopausal status, tumor size, and node status.
The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) [4] aims to understand breast cancer through molecular taxonomy to develop personalized treatments based on tumor genetic profiles. The dataset encompasses a mix of clinical features and genomic data, with a patient cohort from Canada and UK.
The Study to Understand Prognoses Preferences Outcomes and Risks of Treatment (SUPPORT) [26] focuses on critically ill hospitalized patients, conducted in two phases between 1989 and 1997. It includes administrative and clinical follow-up for six months post-inclusion in the study.
The Veteran Administration Lung Cancer Trial (Veterans) [27] dataset focuses on lung cancer patients treated with two different chemotherapy regimens. This dataset is frequently used in simple survival benchmarks due to its small sample size.
The Worcester Heart Attack Study (WHAS) [28] deals with cardiovascular health, tracking 461 patients post-myocardial infarction from 1997 to 2001. It includes biometric parameters and temporal features like hospital stay length and follow-up dates.
@@
AIDS 1151 91.7% 11
Breast Cancer 198 74.2% 80
GBSG 686 56.4% 8
METABRIC 1904 42.1% 9
SUPPORT 9105 31.9% 35
Veterans 137 6.6% 6
WHAS 461 61.8% 16
@@ 4.2 Metrics
We evaluated survival models using concordance index (C-Index) and integrated Brier score (IBS). The C-Index [10] measures the predictive accuracy of survival models by evaluating the proportion of concordant pairs relative to all comparable pairs within a dataset. A pair of subjects $i$ and $j$ is considered comparable if, given $t_{i}<t_{j}$, then $\delta_{i}=1$. A pair of comparable subjects is concordant when the predicted mean time aligns with the actual event times.
The Brier score [11] asses the calibration of probability estimates over time by computing the weighted squared difference between the binary survival indicator of a subject and the predicted survival probability. The Brier score at time $t$ is defined as:
$\text{BS}(t)=\frac{1}{N}\sum\limits_{i=1}^{N}w_{i}(t)(\mathbf{1}(t_{i}>t)-S(t|x_{i}))^{2},$
where $\mathbf{1}(\cdot)$ is an indicator function and $w_{i}(t)$ is a weighting factor that adjusts the censoring bias. This adjustment is the Inverse Probability of Censoring Weighting (IPCW) [10, 29], which assigns weights based on the inverse probability of censoring at a given time as
$w_{i}(t)=\begin{cases}\delta_{i}/G(t_{i})&\text{if }t_{i}\leq t\\ 1/G(t_{i})&\text{if }t_{i}>t\end{cases}.$
where $G(t)$ represents the Kaplan-Meier estimate of the censoring distribution, calculated over the dataset with inverted censoring indicators $\delta$. The overall calibration of a survival model over time can be summarized by integrating the Brier score across the entire study period, yielding the Integrated Brier Score (IBS).
@@ 4.3 Experimental Procedure
To ensure a robust evaluation, each dataset is divided into 10 seeded splits of equal size. All subsequent measurements are averaged over 10 independent executions. During each execution, two consecutive splits are defined as validation and test sets, with the remaining eight as the training set, following a standard 10-fold cross-validation approach. This method ensures that each data sample is considered once in the final test evaluation. Data splits are performed with stratification on the event binary variable to maintain a comparable level of event representation across different splits.
@@
In order to validate FPBoost, we compared its performance against several baseline models. For these models, we employ default parameters provided by implementations in Python survival libraries. Specifically, we utilize scikit-survival [12] for RSF, Cox, and CoxBoost, pycox [18] for DeepSurv and DeepHit, and auton-survival [30] for DSM. Following [4], neural network architectures for DeepSurv, DeepHit, and DSM comprise three layers with neuron counts of 3, 5, and 3 times the number of features, respectively.
@@
For FPBoost, we conduct a hyperparameter search for each dataset, selecting the model with the highest mean validation score, defined as C-Index -- IBS. The search, performed using ray-tune [31], explores a space including the number of Weibull heads ($\{0,\dots,8\}$), the number of LogLogistic heads ($\{0,\dots,8\}$), the number of gradient-boosted trees per parameter ($\{1,\dots,256\}$), the boosting learning rate ($[0.01,1]$), and the ElasticNet loss parameters (${\alpha\in[0,1]}$ and ${\gamma\in[0,1]}$). To prevent fully parametric distributions to be affected by different time scales, we normalized the time values for each dataset between 0 and 1. The source code for these experiments is available at https://github.com/archettialberto/fpboost. The FPBoost class implementation is fully compatible with scikit-survival, facilitating seamless integration into existing codebases.
5 Results and Discussion
@@
Table 3: Test C-Index ($\uparrow$) and 95% confidence interval of each model and dataset averaged across 10 seeded folds. For improved readability, all values reported are scaled up by a factor of 100. For each column, the best result is bold and the second best is underlined.
@@
Model AIDS Breast Cancer GBSG METABRIC SUPPORT Veterans WHAS
RSF 71.78 $\pm$ 3.95 65.86 $\pm$ 2.83 67.52 $\pm$ 3.59 63.50 $\pm$ 1.92 84.05 $\pm$ 0.59 69.61 $\pm$ 5.89 76.40 $\pm$ 3.53
Cox 73.01 $\pm$ 2.70 55.68 $\pm$ 6.21 67.97 $\pm$ 3.74 63.44 $\pm$ 1.12 78.60 $\pm$ 0.86 70.63 $\pm$ 5.67 76.17 $\pm$ 2.90
CoxBoost 70.31 $\pm$ 3.09 64.87 $\pm$ 6.74 67.75 $\pm$ 3.92 63.67 $\pm$ 1.31 85.42 $\pm$ 0.56 65.77 $\pm$ 6.82 75.74 $\pm$ 2.67
DeepSurv 65.03 $\pm$ 6.75 58.14 $\pm$ 8.14 57.49 $\pm$ 2.99 56.58 $\pm$ 3.16 85.36 $\pm$ 0.51 56.57 $\pm$ 7.41 72.28 $\pm$ 3.66
DSM 69.72 $\pm$ 5.47 67.93 $\pm$ 5.68 66.73 $\pm$ 3.62 11.51 $\pm$ 14.3 83.46 $\pm$ 0.69 40.90 $\pm$ 6.08 74.67 $\pm$ 2.25
DeepHit 70.90 $\pm$ 4.29 52.72 $\pm$ 8.97 53.29 $\pm$ 4.49 53.41 $\pm$ 2.18 83.18 $\pm$ 0.67 61.06 $\pm$ 4.76 67.20 $\pm$ 4.85
FPBoost 72.60 $\pm$ 3.16 68.13 $\pm$ 6.35 69.75 $\pm$ 3.51 64.23 $\pm$ 1.39 85.44 $\pm$ 0.56 70.68 $\pm$ 6.64 77.40 $\pm$ 2.88
@@
Table 4: Test IBS ($\downarrow$) and 95% confidence interval of each model and dataset averaged across 10 seeded folds. For improved readability, all values reported are scaled up by a factor of 100. For each column, the best result is bold and the second best is underlined.
@@
Model AIDS Breast Cancer GBSG METABRIC SUPPORT Veterans WHAS
RSF 6.23 $\pm$ 0.42 17.62 $\pm$ 3.34 19.06 $\pm$ 1.14 19.73 $\pm$ 0.89 12.76 $\pm$ 0.60 13.85 $\pm$ 2.25 16.35 $\pm$ 1.39
Cox 6.20 $\pm$ 0.44 36.28 $\pm$ 9.66 18.67 $\pm$ 1.06 19.21 $\pm$ 0.66 13.60 $\pm$ 0.59 14.42 $\pm$ 2.03 15.90 $\pm$ 1.40
CoxBoost 6.32 $\pm$ 0.40 22.40 $\pm$ 8.41 19.03 $\pm$ 1.45 19.64 $\pm$ 0.90 12.97 $\pm$ 0.60 15.29 $\pm$ 2.71 17.27 $\pm$ 1.86
DeepSurv 6.23 $\pm$ 0.36 20.25 $\pm$ 6.38 21.01 $\pm$ 1.24 21.26 $\pm$ 0.89 13.15 $\pm$ 0.59 16.95 $\pm$ 2.58 19.69 $\pm$ 1.01
DSM 6.26 $\pm$ 0.38 17.20 $\pm$ 1.82 22.20 $\pm$ 0.99 24.52 $\pm$ 0.60 19.53 $\pm$ 0.67 23.03 $\pm$ 1.17 21.92 $\pm$ 0.51
DeepHit 7.84 $\pm$ 0.85 20.49 $\pm$ 4.48 22.82 $\pm$ 1.67 22.25 $\pm$ 0.85 15.65 $\pm$ 0.44 32.51 $\pm$ 3.17 23.09 $\pm$ 1.35
FPBoost 6.12 $\pm$ 0.37 17.04 $\pm$ 2.95 18.50 $\pm$ 0.94 19.09 $\pm$ 0.69 12.53 $\pm$ 0.61 14.10 $\pm$ 2.20 17.01 $\pm$ 1.00
Table 5: C-Index and IBS results averaged across all datasets. For improved readability, all values reported are scaled up by a factor of 100. The best result is bold and the second best is underlined.
Model C-Index ($\uparrow$) IBS ($\downarrow$)
RSF 71.24 $\pm$ 3.19 15.09 $\pm$ 1.43
Cox 69.36 $\pm$ 3.31 17.75 $\pm$ 2.26
CoxBoost 70.50 $\pm$ 3.59 16.13 $\pm$ 2.33
DeepSurv 64.49 $\pm$ 4.66 16.93 $\pm$ 1.86
DSM 72.50 $\pm$ 3.54 17.42 $\pm$ 0.87
DeepHit 63.10 $\pm$ 4.32 20.66 $\pm$ 1.83
FPBoost 72.60 $\pm$ 3.50 14.91 $\pm$ 1.25
@@
This section presents and analyzes the empirical evaluation of FPBoost against state-of-the-art survival models described in the Survival Models Section, specifically RSF, Cox, CoxBoost, DeepSurv, DSM, and DeepHit. Tables 3 and 4 report the performance of each model according to the C-Index and IBS metrics, respectively. Additionally, Table 5 summarizes the mean performance across all datasets. For improved readability, all results and metric reports are scaled up by a factor of 100.
@@
The C-Index results in Table 3 demonstrate FPBoost’s competitive performance across all datasets, outperforming other models in all cases except AIDS, where it is marginally surpassed by the Cox model. On average, FPBoost improves the C-Index by 4.07 points across all datasets and models, representing an approximately 8% improvement given that a random guessing model has a C-Index of 50 [2]. When compared specifically to proportional models (Cox, CoxBoost, and DeepSurv), FPBoost’s average improvement is 4.49, highlighting the potential limitations of the proportional hazard assumption in capturing complex data patterns. The performance gap widens further when comparing FPBoost to neural network-based models (DeepSurv, DSM, and DeepHit), with an average improvement of 5.90. This suggests that neural networks, despite their capacity to learn complex non-linear patterns, may require more sophisticated tuning to prevent overfitting and ensure strong generalization. Notably, the average improvement of FPBoost over tree-based methods (RSF and CoxBoost) is smaller at 1.73, indicating that tree-based methods with minimal assumptions may be better suited for tabular survival data compared to neural networks.
@@
Table 4 demonstrates FPBoost’s strong calibration performance according to the IBS metric, corroborating the trends analyzed on C-Index. Here, FPBoost outperforms all baselines except in the Veterans and WHAS datasets, where it ranks as the second-best model in the former. The average IBS improvement of FPBoost across all datasets is 2.42, again representing approximately a 10% improvement, given that a random guessing model has an IBS of 25 [2]. Consistently with the C-Index results, the IBS improvement relative to proportional hazard models is 2.03, increasing to 3.43 for neural network-based models, while the difference with tree-based models is smaller at 0.70.
@@
In summary, the empirical results demonstrate FPBoost’s competitive performance against various state-of-the-art survival models, both tree-based and neural network-based. Performance improvements are evident in terms of both concordance and calibration. These findings suggest that FPBoost’s tree-based nature, combined with direct optimization of the survival likelihood, represents a promising approach for developing complex, competitive, and adaptable survival models.

### v2

@@ 3.2 Loss Function and Training Procedure
where $\alpha\geq 0$ weighs the regularization contribution to the final loss and $\gamma\in[0,1]$ controls the ratio between the $L1$ and $L2$ penalties.
Training is performed using a standard gradient boosting algorithm. Specifically, for each parameter from $\Theta$, being it a scale, shape, or weight, an empty list of trees is initialized. Then, at iteration $m$, each of these lists is populated by a new tree $\tau_{j}^{(m)}(\mathbf{x})$ fitted on the negative gradient of the loss function, called pseudo-residual. Parameter estimation occurs by summing the estimations of all trees belonging to each list, weighted by a learning rate $\lambda>0$ (Friedman, 2001). This way, each iteration produces a set of trees which contribute to loss minimization and consequently to parameter refinement, in a standard gradient boosting fashion. The pseudocode of FPBoost is provided in Algorithm 1.
@@
0:  Training data $\{(\mathbf{x}_{i},\delta_{i},t_{i})\}_{i=1}^{N}$, number of heads $J$, distribution types for each head (Weibull or LogLogistic), number of iterations $M$, learning rate $\lambda$, regularization parameters $\alpha$ and $\gamma$, weight activation function $\phi(\cdot)$
1:  Initialize $3J$ parameter models with random values: ${\bigl{\{}F_{\eta_{j}}^{(0)}(\mathbf{x}),\;F_{k_{j}}^{(0)}(\mathbf{x}),\;F_{w_{j}}^{(0)}(\mathbf{x})\bigr{\}}_{j=1}^{J}}$
@@
3:     for $j=1$ to $J$ do
4:        $\eta_{j}=\mathrm{ReLU}\bigl{(}F_{\eta_{j}}^{(m)}(\mathbf{x})\bigr{)}$
5:        $k_{j}=\mathrm{ReLU}\bigl{(}F_{k_{j}}^{(m)}(\mathbf{x})\bigr{)}$
6:        $w_{j}=\phi\bigl{(}F_{w_{j}}^{(m)}(\mathbf{x})\bigr{)}$
7:     end for
8:     Collect parameters as $\Theta=\{\eta_{j},k_{j},w_{j}\}_{j=1}^{J}$
9:     $L_{\mathrm{lik}}=-\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\delta_{i}\log\bigl{(}h(t_{i}\mid\Theta_{i})\bigr{)}-H(t_{i}\mid\Theta_{i})\Bigr{]}$
10:     $L_{\mathrm{reg}}=\alpha\Bigl{(}\gamma\|\Theta\|_{1}+(1-\gamma)\|\Theta\|_{2}^{2}\Bigr{)}$
11:     $L\;=\;L_{\mathrm{lik}}\;+\;L_{\mathrm{reg}}$
12:     for each parameter model $F_{\cdot_{j}}^{(m)}(\mathbf{x})$ do
13:        Compute pseudo-residuals ${r_{\cdot_{j}}=-\frac{\partial L\bigl{(}\mathbf{x},\delta,t,\Theta\bigr{)}}{\partial F_{\cdot_{j}}^{(m)}(\mathbf{x})}}$
14:        Fit a regression tree $\tau_{\cdot_{j}}^{(m)}$ to $r_{\cdot_{j}}$
15:        Update the parameter model: $F_{\cdot_{j}}^{(m+1)}(\mathbf{x})=F_{\cdot_{j}}^{(m)}(\mathbf{x})+\lambda\;\tau_{\cdot_{j}}^{(m)}(\mathbf{x})$
16:     end for
17:  end for
18:  Return parameter models $\bigl{\{}F_{\eta_{j}}^{(M)},F_{k_{j}}^{(M)},F_{w_{j}}^{(M)}\bigr{\}}_{j=1}^{J}$
@@ 4.1 Datasets
To ensure fair evaluation and consistency with similar studies, we selected datasets from well-known benchmarks in survival analysis, covering different conditions like breast cancer, lung cancer, AIDS, and cardiovascular diseases. Specifically, the AIDS (Hosmer et al., 2008), Breast Cancer (Desmedt et al., 2007), FLCHAIN (Dispenzieri et al., 2012), GBSG2 (Schumacher et al., 1994) and Veterans (Kalbfleisch & Prentice, 2011) datasets are provided by the scikit-survival (Pölsterl, 2020) Python library. The METABRIC (Katzman et al., 2018) and WHAS (Hosmer et al., 2008) datasets, instead, are available in the DeepSurv repository (Katzman et al., 2018) with a predefined test set. Lastly, SUPPORT2 (Knaus et al., 1995) is provided by SurvSet (Drysdale, 2022). Details on data collection and content overview are provided in Appendix A.2. Table 2 collects the summary statistics of these datasets.
@@
AIDS 1151 91.66% 11
Breast Cancer 198 74.24% 80
FLCHAIN 7874 72.45% 9
GBSG2 686 56.41% 8
METABRIC 1904 42.07% 9
SUPPORT2 9105 31.89% 35
Veterans 137 6.57% 6
WHAS 1638 57.88% 6
@@ 4.2 Metrics
We evaluated survival models using the concordance index (C-Index) and the integrated Brier score (IBS). The C-Index (Uno et al., 2011) measures the predictive accuracy of survival models by evaluating the proportion of concordant pairs relative to all comparable pairs within a dataset. A pair of subjects $i$ and $j$ is considered comparable if, given $t_{i}<t_{j}$, then $\delta_{i}=1$. A pair of comparable subjects is concordant when the predicted mean time aligns with the actual event times.
The Brier score (Graf et al., 1999) assesses the calibration of probability estimates over time by computing the weighted squared difference between the binary survival indicator of a subject and the predicted survival probability. The Brier score at time $t$ is defined as:
$\text{BS}(t)=\frac{1}{N}\sum\limits_{i=1}^{N}w_{i}(t)(\mathbf{1}(t_{i}>t)-S(t|\mathbf{x}_{i}))^{2},$
where $\mathbf{1}(\cdot)$ is an indicator function and $w_{i}(t)$ is a weighting factor that adjusts the censoring bias. This adjustment is the Inverse Probability of Censoring Weighting (IPCW) (Robins & Rotnitzky, 1992; Uno et al., 2011), which assigns weights based on the inverse probability of censoring at a given time $t$. The overall calibration of a survival model over time is summarized by integrating the Brier score across the entire study period, yielding the Integrated Brier Score (IBS).
@@ 4.3 Experimental Procedure
Each dataset is split into a training and test set, with the latter accounting for 20% of the total samples. For datasets coming from the DeepSurv repository, we employ the provided train-test split while for the others we apply a seeded stratified split on the censoring variable. To ensure a robust evaluation, each experiment is run 30 times with different seeds and all subsequent measurements are averaged over all executions. During each of these executions, the training set is further divided into training and validation, to allow for hyperparameter tuning and early stopping. Before training, standard normalization and one-hot encoding are applied to numerical and categorical features, respectively.
@@
In order to validate FPBoost, we compared its performance against several baseline models. For these models, we employ default parameters provided by implementations in Python libraries. Specifically, we utilize scikit-survival for RSF, Cox, and CoxBoost, pycox (Kvamme et al., 2019) for DeepSurv and DeepHit, xgboost (Chen et al., 2022) for XGBoost, and auton-survival (Nagpal et al., 2022) for DSM. Following (Katzman et al., 2018), neural network architectures for DeepSurv, DeepHit, and DSM comprise three layers with neuron counts of 3, 5, and 3 times the number of features, respectively.
@@
For FPBoost, we conduct a hyperparameter search for each dataset, selecting the model with the highest mean C-Index. The search, performed using random search, explores a space including the number of Weibull heads ($\{0,\dots,32\}$), the number of LogLogistic heads ($\{0,\dots,32\}$), the number of gradient-boosted trees per parameter ($\{1,\dots,512\}$), the maximum tree depth ($\{1,\dots,6\}$), the weights activation functions (ReLU, sigmoid, softmax, hyperbolic tangent, or identity), the boosting learning rate ($[0.01,1]$), the and ElasticNet loss parameters (${\alpha\in[0,1]}$ and ${\gamma\in[0,1]}$). To prevent fully parametric distributions to be affected by different time scales, we normalized the time values for each dataset between 0 and 1. The best hyperparameters for each dataset are provided in Appendix A.3.
The source code for these experiments is available at [Anonymized Link]. The FPBoost class implementation is fully compatible with scikit-survival, facilitating its inclusion into existing codebases.
5 Results
@@
Table 3: Test C-Index ($\uparrow$) and 95% confidence interval for each model and dataset, averaged across 30 seeded splits (same test, different train and validation sets). To enhance readability, all values are scaled by a factor of 100. Best results are highlighted in bold, while the second best are underlined.
@@
Model AIDS Breast FLCHAIN GBSG2 METABRIC SUPPORT2 Veterans WHAS
Cox $77.8\pm 1.9$ $63.4\pm 1.9$ $93.7\pm 0.0$ $69.3\pm 0.3$ $63.2\pm 0.1$ $82.7\pm 0.0$ $75.6\pm 1.0$ $81.7\pm 0.1$
CoxBoost $76.1\pm 0.9$ $60.0\pm 2.6$ $93.7\pm 0.0$ $68.9\pm 0.6$ $63.2\pm 0.2$ $83.4\pm 0.0$ $72.1\pm 1.9$ $85.1\pm 0.1$
XGBoost $53.2\pm 1.5$ $57.1\pm 2.7$ $88.9\pm 0.1$ $63.4\pm 0.9$ $61.4\pm 0.5$ $56.4\pm 0.7$ $70.9\pm 1.5$ $83.4\pm 0.3$
RSF $\mathbf{80.1\pm 0.8}$ $58.5\pm 1.6$ $93.7\pm 0.0$ $68.6\pm 0.4$ $61.6\pm 0.2$ $84.2\pm 0.1$ $75.8\pm 1.0$ $85.8\pm 0.1$
DeepSurv $70.7\pm 3.0$ $64.9\pm 1.7$ $93.6\pm 0.0$ $69.5\pm 0.4$ $63.4\pm 0.2$ $82.6\pm 0.1$ $\mathbf{76.7\pm 1.0}$ $83.7\pm 0.1$
DeepHit $78.4\pm 0.9$ $64.6\pm 2.4$ $93.5\pm 0.0$ $65.5\pm 1.0$ $61.7\pm 0.3$ $82.2\pm 0.1$ $72.0\pm 1.4$ $82.6\pm 0.2$
DSM $76.8\pm 1.0$ $66.2\pm 0.8$ $50.0\pm 0.0$ $49.9\pm 0.3$ $61.3\pm 0.1$ $83.6\pm 0.3$ $65.4\pm 0.2$ $69.7\pm 0.6$
FPBoost $78.1\pm 0.7$ $\mathbf{66.6\pm 3.2}$ $\mathbf{93.8\pm 0.0}$ $\mathbf{69.7\pm 0.4}$ $\mathbf{64.0\pm 0.1}$ $\mathbf{84.3\pm 0.4}$ $74.8\pm 1.3$ $\mathbf{89.0\pm 0.3}$
@@
Table 4: Test IBS ($\downarrow$) and 95% confidence interval for each model and dataset, averaged across 30 seeded splits (same test, different train and validation sets). To enhance readability, all values are scaled by a factor of 100. Best results are highlighted in bold, while the second best are underlined.
@@
Model AIDS Breast FLCHAIN GBSG2 METABRIC SUPPORT2 Veterans WHAS
Cox $\mathbf{5.8\pm 0.0}$ $22.2\pm 1.2$ $\mathbf{4.6\pm 0.0}$ $17.7\pm 0.1$ $19.9\pm 0.0$ $13.2\pm 0.0$ $13.4\pm 0.2$ $14.0\pm 0.0$
CoxBoost $6.2\pm 0.1$ $19.8\pm 0.9$ $4.7\pm 0.0$ $17.3\pm 0.2$ $20.9\pm 0.1$ $12.6\pm 0.0$ $14.7\pm 0.7$ $11.9\pm 0.1$
XGBoost $8.5\pm 0.2$ $22.8\pm 0.9$ $12.1\pm 0.1$ $26.9\pm 0.7$ $24.3\pm 0.4$ $60.6\pm 0.6$ $47.4\pm 1.7$ $18.7\pm 0.4$
RSF $5.8\pm 0.0$ $18.0\pm 0.3$ $4.7\pm 0.0$ $17.7\pm 0.2$ $21.0\pm 0.1$ $\mathbf{12.3\pm 0.0}$ $\mathbf{12.7\pm 0.2}$ $8.5\pm 0.1$
DeepSurv $6.2\pm 0.1$ $26.4\pm 1.1$ $4.7\pm 0.0$ $17.5\pm 0.1$ $20.4\pm 0.1$ $14.6\pm 0.1$ $14.4\pm 0.2$ $12.2\pm 0.1$
DeepHit $5.8\pm 0.0$ $23.9\pm 1.3$ $6.3\pm 0.1$ $21.4\pm 0.1$ $22.7\pm 0.1$ $14.7\pm 0.1$ $29.4\pm 0.9$ $17.2\pm 0.1$
DSM $6.2\pm 0.0$ $17.8\pm 0.0$ $13.9\pm 0.0$ $21.6\pm 0.0$ $23.7\pm 0.0$ $19.4\pm 0.3$ $23.0\pm 0.1$ $20.5\pm 0.0$
FPBoost $6.0\pm 0.0$ $\mathbf{17.1\pm 0.9}$ $4.7\pm 0.0$ $\mathbf{17.1\pm 0.2}$ $\mathbf{19.8\pm 0.0}$ $12.5\pm 0.1$ $12.8\pm 0.3$ $\mathbf{8.4\pm 0.3}$
@@
This section presents and analyzes the empirical evaluation of FPBoost against classical and state-of-the-art survival models described in Section 2, specifically RSF, Cox, CoxBoost, DeepSurv, DSM, and DeepHit. Tables 3 and 4 report the performance of each model according to the C-Index and IBS metrics, respectively. Additional results and metric summaries across model types are provided in Appendix A.4. For improved readability, all results and metric reports are scaled up by a factor of 100.
@@
The C-Index results in Table 3 demonstrate FPBoost’s competitive performance across all datasets, outperforming other models in all cases except for AIDS and Veteran datasets, where it is marginally surpassed by RSF and DeepSurv. Averaging across all the datasets, FPBoost improves the C-Index against the baseline score by 4.6 points. Since, by definition, a random guessing model has a C-Index of 50 and the C-Index metric is evaluated in a $[50,100]$ range, the performance gain of FPBoost on said metric approximately corresponds to a 9% improvement. When compared specifically to semi-parametric models (Cox, CoxBoost, and DeepSurv), FPBoost’s average improvement is 4.1, highlighting the potential limitations of the proportional hazard assumption in capturing complex data patterns. However, the performance gap in favor of FPBoost becomes more pronounced when compared to neural network-based models (DeepSurv, DSM, and DeepHit), showing an average improvement of 5.5. This suggests that neural networks, despite their capacity to learn complex non-linear patterns, may require more sophisticated tuning to prevent overfitting and ensure strong generalization. In such cases, simplifying assumptions can be beneficial in introducing bias, as suggested by DeepSurv surpassing both DeepHit and DSM. Notably, the average improvement of FPBoost over RSF is smaller at 1.5, indicating that non-parametric algorithms with minimal assumptions may be better suited for tree-based learners compared to neural networks.
@@
Table 4 demonstrates the calibration performance of FPBoost according to the IBS metric, corroborating the trends analyzed on C-Index. Here, FPBoost always ranks first or second, with the exception of the AIDS dataset. Averaging across all datasets, the improvement in IBS of FPBoost is 2.8. Since, by definition, a random guessing model has an IBS of 25 and the IBS metric is evaluated in a $[0,25]$ range, the raw score improvement translates to an approximate 11% improvement. We opted for excluding XGBoost from these calculations, given its outlier performance with respect to the alternatives. Consistently with the C-Index results, the IBS improvement relative to proportional hazard models is 1.7, increasing to 4.5 for neural network-based models, while the difference with tree-based models is smaller at 0.7. These findings are in line with C-Index, indicating that model performance on survival concordance is reflected also on probability calibration.
@@
In summary, these empirical results showcase the competitive performance of FPBoost against various classical and state-of-the-art models, both tree-based and neural network-based. Performance improvements are evident in terms of both concordance and calibration with respect to neural alternatives. These findings suggest that tree-based nature of FPBoost, combined with direct optimization of the survival likelihood, represents a promising approach for developing more complex, competitive, and adaptable survival models.
5.1 Discussion and Future Work
While the idea of leveraging a mixture of parametric functions (Nagpal et al., 2021) has been previously explored in the survival literature, as well as ensemble learning (Ishwaran et al., 2008; Archetti et al., 2023), FPBoost introduces several innovations.
Firstly, the weighted sum is applied directly to the hazard function, unlike previous works that applied it to the survival function. The advantage of this formulation is twofold. On the one hand, summation on Weibull hazards guarantees universal approximation, provided enough heads. On the other hand, a direct comparison with DSM suggests that learning parameters directly in hazard-space rather than explicitly weighting survival distributions can be beneficial for a more effective training.
Secondly, FPBoost directly maximizes the survival likelihood, without relying on simplified custom loss functions such as the partial likelihood of proportional models or discrete loss functions of neural-network-based models. This is possible due to the assumption that the global hazard function is a composition of differentiable parametric hazard functions. This aspect, combined with the tree-based nature of the algorithm, contributed the most to the empirical results obtained. In fact, ensembles of decision tree have been historically extremely effective in dealing with tabular data so far, even against neural networks and deep learning (Grinsztajn et al., 2022).
Future work could explore the theoretical bounds and practical limits of the approximation capabilities of FPBoost, building upon the results presented in this study to improve its mathematical grounding. Another aspect worth investigating is the application of FPBoost to competing risks scenarios, for example by delegating separate sets of heads to different events of interest, as in DeepHit. On top of that, investigating model performance on larger, more diverse and multimodal datasets beyond the healthcare context could further validate its practical utility. Finally, its inclusion in federated learning scenarios can be beneficial for applications where data scarcity and privacy hinder the results and applicability of existing models.